### v2

# OmniSync: Towards Universal Lip Synchronization via Diffusion Transformers

Abstract
Lip synchronization is the task of aligning a speaker’s lip movements in video with corresponding speech audio, and it is essential for creating realistic, expressive video content. However, existing methods often rely on reference frames and masked-frame inpainting, which limit their robustness to identity consistency, pose variations, facial occlusions, and stylized content. In addition, since audio signals provide weaker conditioning than visual cues, lip shape leakage from the original video will affect lip sync quality. In this paper, we present OmniSync, a universal lip synchronization framework for diverse visual scenarios. Our approach introduces a mask-free training paradigm using Diffusion Transformer models for direct frame editing without explicit masks, enabling unlimited-duration inference while maintaining natural facial dynamics and preserving character identity. During inference, we propose a flow-matching-based progressive noise initialization to ensure pose and identity consistency, while allowing precise mouth-region editing. To address the weak conditioning signal of audio, we develop a Dynamic Spatiotemporal Classifier-Free Guidance (DS-CFG) mechanism that adaptively adjusts guidance strength over time and space. We also establish the AIGC-LipSync Benchmark, the first evaluation suite for lip synchronization in diverse AI-generated videos. Extensive experiments demonstrate that OmniSync significantly outperforms prior methods in both visual quality and lip sync accuracy, achieving superior results in both real-world and AI-generated videos.
Figure 1: OmniSync demonstrates universal lip synchronization capabilities, effectively handling facial occlusion, while maintaining visual consistency and generating accurate lip movements.
1 Introduction
Lip synchronization, matching mouth movements with speech audio, is essential for creating compelling visual content across film dubbing zhang2023dinet , digital avatars peng2023emotalk ; peng2023selftalk ; zhou2024meta , and teleconferencing meng2024comprehensive ; peng2025dualtalk . With the rise of AI-generated content, this technology has evolved from a specialized technique to a fundamental aspect of the video generation landscape yan2021videogpt ; singer2022make ; liu2024evalcrafter . Despite significant advances in text-to-video (T2V) models chen2023videocrafter1 ; blattmann2023stable ; yang2024cogvideox ; kong2024hunyuanvideo ; wang2025wan creating increasingly realistic footage, achieving precise and natural lip synchronization remains an unsolved challenge.
Traditional lip synchronization approaches rely heavily on reference frames combined with masked-frame inpainting prajwal2020lip ; zhang2023dinet ; guan2023stylesync ; guan2024resyncer . This methodology extracts appearance information from reference frames to inpaint masked regions in target frames—a process that introduces several critical limitations. These methods struggle with head pose variations, identity preservation, and artifact elimination, especially when target poses differ significantly from references peng2024synctalk ; peng2025synctalk++ .
Furthermore, the dependence on explicit masks cannot fully prevent unwanted lip shape leakage, compromising synchronization quality and restricting applicability across diverse visual representations bigata2025keysync . The challenges intensify in the context of audio-driven generation. Unlike strong visual cues, audio signals provide relatively weak conditioning, making precise lip synchronization difficult wang2024v . Additionally, existing methods rely on face detection and alignment bulat2017far techniques that break down when applied to stylized characters and non-human entities, precisely the diverse content that modern text-to-video models excel at generating.
This technical gap is compounded by the absence of standardized evaluation frameworks for lip sync in stylized videos. Current benchmarks zhang2021flow ; wang2020mead focus almost exclusively on photorealistic human faces in controlled settings, failing to capture the visual diversity inherent in AI-generated videos.
To address these challenges, we introduce OmniSync, a universal lip synchronization framework designed for diverse videos. Our approach eliminates reliance on reference frames and explicit masks through a diffusion-based direct video editing paradigm. In addition, we establish AIGC-LipSync Benchmark, the first comprehensive evaluation framework for lip synchronization across diverse AIGC contexts. OmniSync’s technical approach is built upon three key innovations:
First, we implement a mask-free training paradigm using Diffusion Transformers (DiT) peebles2023scalable for direct cross-frame editing. Our model learns a mapping function $(V_{cd},A_{ab})\mapsto V_{ab}$, where $V$ represents video frames and $A$ represents audio. The indices $(a:b,c:d)$ represent different segments sampled from the same video. The model modifies only speech-relevant regions according to target audio without requiring explicit masks or references. This approach enables unlimited-duration inference while maintaining natural facial dynamics and preserving character identity.
Second, we introduce a flow-matching-based progressive noise initialization strategy during inference. Rather than beginning with random noise tian2024emo , we inject controlled noise into original frames using Flow Matching lipman2022flow , then execute only the final denoising steps. This approach maintains spatial consistency between source and generated frames while allowing sufficient flexibility for precise mouth region modifications, effectively mitigating pose inconsistency and identity drift.
Third, we develop a dynamic spatiotemporal Classifier-Free Guidance (CFG) framework ho2022classifier that provides fine-grained control over the generation process. By adaptively adjusting guidance strength across both temporal and spatial dimensions: temporally reducing guidance strength as denoising progresses, and spatially applying Gaussian-weighted control centered on mouth-relevant regions. This balanced approach ensures precise lip synchronization without disturbing unrelated areas.
Our contributions can be summarized as follows:
  • $\bullet$
    A universal lip synchronization framework that eliminates reliance on reference frames and explicit masks, enabling accurate speech synchronization across diverse visual representations.
  • $\bullet$
    A flow-matching-based progressive noise initialization strategy during inference, effectively stabilizing the early denoising process and mitigating pose inconsistency and identity drift.
  • $\bullet$
    A dynamic spatiotemporal CFG framework that provides fine-grained control over audio influence, addressing the weak signal problem in audio-driven generation.
  • $\bullet$
    A comprehensive AIGC-LipSync Benchmark for evaluating lip synchronization in AI-generated content, including stylized characters and non-human entities.
2 Related Work
2.1 Audio-driven Lip Synchronization
GAN-based Lip Synchronization. Traditional GAN-based goodfellow2020generative methods prajwal2020lip ; wang2023seeing ; cheng2022videoretalking ; ma2023styletalk ; gupta2023towards have established important foundations in lip synchronization. Wav2Lip prajwal2020lip pioneered the use of pretrained SyncNet to supervise generator training, setting a benchmark for subsequent research. DINet zhang2023dinet enhanced synchronization quality by performing spatial deformation on reference image feature maps, better preserving high-frequency details. IP-LAP zhong2023identity introduced a two-stage approach that first infers landmarks from audio before rendering them into facial images. ReSyncer guan2024resyncer incorporated 3D mesh priors for facial motion, effectively reducing artifacts.
Diffusion-based Lip Synchronization. Recent advances in diffusion models mukhopadhyay2024diff2lip ; zhang2024musetalk ; li2024latentsync ; ma2025sayanything have enabled significant progress in audio-driven lip synchronization. LatentSync li2024latentsync represents an end-to-end framework based on audio-conditioned latent diffusion models without intermediate motion representation. SayAnything ma2025sayanything employs a denoising UNet architecture that processes video latents with audio conditioning. MuseTalk zhang2024musetalk proposes a novel sampling strategy that selects reference images with head poses closely matching the target.
However, these methods still rely on reference frames combined with masked-frame inpainting, leading to head pose limitations, identity preservation issues, and blurry edge generation. Our OmniSync framework addresses these limitations through a mask-free training paradigm that enables application across diverse visual representations.
2.2 Audio-driven Portrait Animation
Audio-driven portrait animation tian2024emo ; xu2024hallo ; jiang2024loopy ; chen2025echomimic ; ji2024sonic differs fundamentally from lip sync. Portrait animation wei2024aniportrait ; cui2024hallo2 follows an image-to-video framework without constraints on head poses or facial expressions, eliminating the need to integrate generated content back into original video. This approach is unsuitable for post-generation lip synchronization in video generation pipelines. In contrast, lip synchronization prajwal2020lip ; li2024latentsync operates within a video-to-video framework, modifying only lip movements while maintaining compatibility with existing footage. This represents a more constrained task, requiring precise modification of lip regions while preserving surrounding facial features.
Recent models like OmniHuman-1 lin2025omnihuman and Mocha wei2025mocha use audio directly as a conditioning signal for image-to-video or text-to-video frameworks. However, due to limitations in talking head datasets, their generative capabilities don’t match the versatility of advanced video generation models. This gap highlights why specialized lip synchronization for AI-generated videos remains critical.
3 Method
3.1 Overview
In this section, we present OmniSync, a universal lip synchronization framework designed for diverse visual content (Fig. 2). Our approach comprises three key components: 1) a mask-free training paradigm that eliminates dependency on reference frames and explicit masks, 2) a flow-matching-based progressive noise initialization strategy for enhanced inference stability, and 3) dynamic spatiotemporal Classifier-Free Guidance (CFG) that optimizes lip sync while preserving facial details. The following subsections provide comprehensive explanations of each component.
Figure 2: Overview of OmniSync. A mask-free training paradigm employs timestep-dependent sampling to predict the lip-synchronized targets $V_{ab}$. During inference, progressive noise initialization and dynamic spatiotemporal CFG ensure consistent head pose and precise lip synchronization.
3.2 Mask-Free Training Paradigm
Traditional lip synchronization methods prajwal2020lip ; zhang2024musetalk rely on masked-frame inpainting, isolating the mouth region before generating content based on audio input. Despite their prevalence, these approaches produce boundary artifacts and struggle with identity preservation. Crucially, they require explicit face detection and alignment—techniques that fail with stylized characters and non-human entities.
An alternative approach is direct frame editing, which aims to transform frames according to target audio without relying on masks or references. However, this approach requires perfectly paired training data with identical head poses and identity—differing only in lip movements. Such paired data is extremely rare and would severely restrict the model’s generalizability to diverse visual results.
To address these limitations, we leverage the progressive denoising characteristic of diffusion models, introducing a novel training strategy that varies data sampling based on diffusion timesteps. This allows for stable learning without requiring perfectly paired examples. Our goal is to learn a conditional generation process mapping $(V_{cd},A_{ab})\mapsto V_{ab}$ through iterative denoising, where $V$ represents video frames and $A$ represents audio.
We employ Flow Matching lipman2022flow as our training objective. Given an input video segment $V_{cd}$ from frames $c$ to $d$, and a target audio segment $A_{ab}$ from frames $a$ to $b$, our model generates the corresponding video segment $V_{ab}$ via the diffusion process:
$x_{t-1}=\text{DiT}(x_{t},V_{cd},A_{ab},t),$ (1)
where $x_{t}$ represents the noised version of target video $V_{ab}$ at timestep $t$, and DiT denotes our diffusion transformer, which predicts the denoised state at timestep t1t-1.
The CFM loss used for training is defined as:
$\mathcal{L}_{CFM}(\theta)=\mathbb{E}_{t,x_{t},V_{cd},A_{ab},V_{ab}}\left[\|v_{\theta}(x_{t},V_{cd},A_{ab},t)-u_{t}(x_{t}|V_{ab})\|_{2}^{2}\right],$ (2)
where $v_{\theta}(x_{t},V_{cd},A_{ab},t)$ is the learned velocity field predicted by DiT with conditioning on input video $V_{cd}$ and target audio $A_{ab}$, $u_{t}(x_{t}|V_{ab})$ is the conditional velocity field typically defined as $u_{t}(x_{t}|V_{ab})=(V_{ab}-x_{t})/(1-t)$ for the linear interpolation path $x_{t}=(1-t)x_{0}+tV_{ab}$.
Timestep-Dependent Sampling Strategy. A critical insight in our approach is recognizing that the diffusion process can be decomposed into distinct phases, each with different learning requirements. Specifically, early timesteps focus on generating fundamental facial structure, including pose and identity information; middle timesteps primarily generate lip movements driven by audio; while late timesteps refine identity details and textures. To capitalize on this natural progression, we utilize different datasets for distinct timesteps.
For early timesteps (approximately $t\approx T$), responsible for generating overall facial structure, we employ pseudo-paired data from controlled laboratory settings. These samples maintain nearly identical pose information, with variations only in lip movements, providing stable learning signals for structural features and ensuring pose alignment between input and output.
For middle and late timesteps, we transition to more diverse data, sampling from arbitrary videos. During middle timesteps, the model learns to generate lip shapes guided by audio input, whereas in late timesteps (approximately $t\approx 0$), it focuses on refining identity details and ensuring texture consistency. This timestep-dependent training strategy can be formalized as:
$p(V_{cd},V_{ab}|t)=\begin{cases}p_{\text{pseudo-paired}}(V_{cd},V_{ab})&\text{if }t>t_{\text{threshold}},\\ p_{\text{arbitrary}}(V_{cd},V_{ab})&\text{otherwise}.\end{cases}$ (3)
Here, $p_{\text{pseudo-paired}}$ indicates sampling from controlled datasets with minimal pose variations, while $p_{\text{arbitrary}}$ signifies sampling from our diverse collection of videos. The conditional generation process can be expressed mathematically as:
$p_{\theta}(V_{ab}|V_{cd},A_{ab})=\int p_{\theta}(V_{ab}|x_{0})p_{\theta}(x_{0}|V_{cd},A_{ab})dx_{0},$ (4)
where $p_{\theta}(V_{ab}|x_{0})$ represents the mapping from the fully denoised state to the output video, and $p_{\theta}(x_{0}|V_{cd},A_{ab})$ captures the relationship between input conditions and the denoised state. Here, $x_{0}$ refers to the completely denoised latent representation (at timestep $t=0$).
This progressive training approach aligns well with the natural learning progression of diffusion models. By providing appropriate training signals at each stage, we enable stable learning even without perfectly paired data, allowing our model to generalize effectively to diverse real-world scenarios while maintaining identity consistency.
3.3 Progressive Noise Initialization
Standard diffusion-based generation tian2024emo typically begins from random noise (timestep $T$) and progressively denoises toward the final output (timestep 0). However, this approach often results in subtle but noticeable pose misalignments between generated content and original video frames, creating undesirable boundary artifacts and compromising identity preservation.
The fundamental issue lies in error accumulation during the early stages of diffusion. Even minor deviations in early timesteps—when basic facial structure is being formed—can lead to significant misalignments in the final output. This problem is relevant for lip synchronization, where the goal is to modify only speech-relevant regions while maintaining perfect spatial consistency elsewhere. To address this challenge, we introduce a flow-matching-based progressive noise initialization strategy that transforms the traditional diffusion process.
Flow-Matching Noise Initialization. Rather than starting the diffusion process from random noise at timestep $T$, we initialize from original video frames with a controlled level of noise. This simulates an intermediate state in the diffusion trajectory, corresponding to a normalized parameter $\tau$. The initialization is performed by adding this controlled noise to the original video frame:
$x_{\text{init}}=\text{FM}_{\text{add}}(V_{\text{source}},\tau)=(1-\tau)V_{\text{source}}+\tau\epsilon,$ (5)
where $x_{\text{init}}$ is the initial noised state derived using the parameter $\tau$, $V_{\text{source}}$ is the source video frame, and $\epsilon\sim\mathcal{N}(0,I)$ is random noise. Let $t_{\text{start}}$ be the discrete timestep corresponding to this initialization point ($T$ is the total number of diffusion steps, and $\tau\in[0,1]$).
This initialization strategy provides two significant advantages. First, it bypasses the early stages of diffusion (from $T$ down to $t_{\text{start}}$) where general facial structure is formed. This ensures that head pose and global structure are directly inherited from the source frame. Second, it reduces computational requirements by performing denoising only for the remaining steps, from $t_{\text{start}}$ down to 0.
The complete progressive denoising process can be expressed as:
$x_{t}=\begin{cases}x_{\text{init}}&\text{if }t=t_{\text{start}},\\ \text{DiT}(x_{t+1},V_{\text{source}},A_{\text{target}},t+1)&\text{if }t_{\text{start}}>t\geq 0,\end{cases}$ (6)
where $A_{\text{target}}$ is the target audio used to guide the denoising process, and $t$ here represents discrete diffusion timesteps.
This approach effectively creates a two-stage process: (1) initialization using the flow-matching-inspired noise addition (Eq. 5) to reach a state equivalent to timestep $t_{\text{start}}$, and (2) guided denoising from $t_{\text{start}}$ to 0 that focuses on modifying mouth regions according to the target audio while preserving the overall facial structure, identity features, and head pose from the source frame. By skipping the early noisy stages where basic structures form, we maintain spatial consistency while allowing sufficient flexibility for precise mouth region modifications.
3.4 Dynamic Spatiotemporal Classifier-Free Guidance
Audio-driven lip synchronization faces a fundamental challenge: audio signals provide relatively weak conditioning compared to visual cues wang2024v . Standard Classifier-Free Guidance (CFG) ho2022classifier can enhance audio conditioning, but applying uniform guidance across spatial and temporal dimensions creates an unavoidable dilemma: higher guidance scales produce more accurate lip sync but introduce texture artifacts, while lower scales preserve visual fidelity but yield less precise lip movements.
To resolve this tension, we introduce Dynamic Spatiotemporal Classifier-Free Guidance (DS-CFG), a novel approach that provides fine-grained control over the generation process across both spatial and temporal dimensions. Our method applies varying guidance strengths to different regions of the frame and different stages of the diffusion process, achieving an optimal balance between lip synchronization accuracy and overall visual quality.
Spatially-Adaptive Guidance. The key insight for spatial adaptation is that audio information primarily affects the mouth region, while other facial areas should remain largely unchanged. We implement this through a Gaussian-weighted spatial guidance matrix that concentrates guidance strength around speech-relevant regions:
$\mathbf{G}_{\text{spatial}}(x,y)=\omega_{\text{base}}+(\omega_{\text{peak}}-\omega_{\text{base}})\cdot\exp\left(-\frac{(x-x_{m})^{2}+(y-y_{m})^{2}}{2\sigma^{2}}\right)$ (7)
where $(x_{m},y_{m})$ represents the mouth center, $\sigma$ controls the spread of the Gaussian distribution, $\omega_{\text{base}}$ is the baseline guidance strength applied to non-mouth regions, and $\omega_{\text{peak}}$ is the peak strength applied at the mouth center. This spatial adaptation ensures that audio conditions strongly influence lip and surrounding regions while minimally affecting other facial features.
Temporally-Adaptive Guidance. We observe that audio conditioning plays different roles at different stages of the diffusion process. In early diffusion timesteps, strong guidance helps establish correct lip shapes, while in later stages, excessive guidance can disrupt fine texture details. To address this, we implement a temporally decreasing guidance schedule:
$\omega(t)=\omega_{\text{peak}}\cdot\left(\frac{t}{T}\right)^{\gamma}$ (8)
where $t$ is the current diffusion timestep, $T$ is the total number of timesteps, $\omega_{\text{peak}}$ is the maximum guidance scale, and $\gamma$ controls the decay rate, with a value of 1.5. This temporal adaptation ensures strong guidance during early and middle diffusion stages when coarse structures are formed, gradually reducing influence during later stages when fine details and textures are refined.
Unified Dynamic Spatiotemporal CFG. Combining both spatial and temporal adaptations, our DS-CFG approach modifies the standard CFG formulation to:
$\hat{\epsilon}_{\theta}(x_{t},c,t)=\epsilon_{\theta}(x_{t},\emptyset,t)+\mathbf{G}_{\text{spatial}}\cdot\omega(t)\cdot(\epsilon_{\theta}(x_{t},c,t)-\epsilon_{\theta}(x_{t},\emptyset,t))$ (9)
where $\epsilon_{\theta}(x_{t},c,t)$ and $\epsilon_{\theta}(x_{t},\emptyset,t)$ are the noise predictions with and without conditioning, respectively.
Through this DS-CFG, our method achieves precise control over the generation process, effectively addressing the weak audio signal problem in audio-driven generation.
4 Experiments
4.1 Experimental Settings
Datasets. We trained OmniSync using the MEAD dataset wang2020mead and a 400-hour dataset collected from YouTube. MEAD’s controlled laboratory recordings with diverse facial expressions but minimal head movement provided ideal data for training early denoising stages, while the YouTube dataset enhanced generalization across varied real-world conditions for middle and late stages.
To address the limitations of existing benchmarks that focus solely on real-world videos with frontal views and stable lighting, we created the AIGC-LipSync Benchmark. This comprehensive evaluation framework comprises 615 human-centric videos generated by state-of-the-art text-to-video models such as Kling, Dreamina, Wan wang2025wan , and Hunyuan kong2024hunyuanvideo . The benchmark specifically captures challenging visual scenarios such as large facial movements, profile views, variable lighting, occlusions, and stylized characters—conditions that traditional benchmarks fail to address. Details about benchmark construction can be found in the supplementary materials.
Implementation Details. We implement our OmniSync model using the Diffusion Transformer architecture. The model is trained on a combined dataset for 80,000 steps using AdamW optimizer loshchilov2017decoupled with a learning rate of 1e-5. Training is completed in 80 hours using 64 NVIDIA A100 GPUs with a batch size of 64. Audio features are extracted via a pre-trained Whisper encoder, and text conditioning utilizes a T5 encoder. Training employs the timestep‑dependent sampling threshold $t_{\text{threshold}}=850$. The experimental results indicate that excessive thresholds induce significant misalignment while insufficient values will leak the original lip shape. During inference we adopt our flow‑matching‑based progressive noise initialization starting at $\tau=0.92$, followed by 50 denoising steps.
Table 1: Quantitative comparison with previous methods on HDTF Dataset.
HDTF Dataset
Method Full Reference Metrics No Reference Metrics Lip Sync
FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ NIQE $\downarrow$ BRISQUE $\downarrow$ HyperIQA $\uparrow$ LMD $\downarrow$ LSE-C $\uparrow$
Wav2Lip prajwal2020lip 14.912 543.340 0.852 6.495 53.372 45.822 10.007 7.630
VideoReTalking cheng2022videoretalking 11.868 379.518 0.786 6.333 50.722 48.476 8.848 7.180
TalkLip wang2023seeing 16.680 691.518 0.843 6.377 52.109 44.393 15.954 5.880
IP-LAP zhong2023identity 9.512 325.691 0.809 6.533 54.402 50.086 7.695 7.260
Diff2Lip mukhopadhyay2024diff2lip 12.079 461.341 0.869 6.261 49.361 48.869 18.986 7.140
MuseTalk zhang2024musetalk 8.759 231.418 0.862 5.824 46.003 55.397 8.701 6.890
LatentSync li2024latentsync 8.518 216.899 0.859 6.270 50.861 53.208 17.344 8.050
Ours 7.855 199.627 0.875 5.481 37.917 56.356 7.097 7.309
Table 2: Quantitative comparison with previous methods on AIGC-LipSync Benchmark.
AIGC-LipSync Benchmark
Method Full Reference Metrics No Reference Metrics Generation Success Rate
FID $\downarrow$ FVD $\downarrow$ CSIM $\uparrow$ NIQE $\downarrow$ BRISQUE $\downarrow$ HyperIQA $\uparrow$
All Videos $\uparrow$
Stylized
Characters $\uparrow$
Wav2Lip prajwal2020lip 22.989 562.245 0.727 5.392 42.816 50.511 71.38% 26.67%
VideoReTalking cheng2022videoretalking 20.439 329.460 0.669 5.947 45.047 48.645 48.78% 7.78%
TalkLip wang2023seeing 31.180 619.179 0.754 5.239 41.692 50.608 52.36% 34.44%
IP-LAP zhong2023identity 14.686 247.402 0.796 5.546 45.153 53.174 45.53% 6.67%
Diff2Lip mukhopadhyay2024diff2lip 23.542 403.149 0.692 5.440 42.442 50.335 74.63% 36.67%
MuseTalk zhang2024musetalk 17.668 297.621 0.667 4.935 36.017 58.334 92.20% 67.78%
LatentSync li2024latentsync 15.374 263.111 0.751 5.342 41.917 54.648 74.96% 35.56%
Ours 10.681 211.350 0.808 4.588 25.485 61.906 97.40% 87.78%
4.2 Quantitative Evaluation
We evaluate OmniSync against state-of-the-art methods including Wav2Lip prajwal2020lip , VideoReTalking cheng2022videoretalking , TalkLip wang2023seeing , IP-LAP zhong2023identity , Diff2Lip mukhopadhyay2024diff2lip , MuseTalk zhang2024musetalk , and LatentSync li2024latentsync using a comprehensive suite of metrics. For visual quality assessment, we employ FID (Fréchet Inception Distance) to measure frame-level fidelity, FVD (Fréchet Video Distance) for temporal consistency, and CSIM (Cosine Similarity) to quantify identity preservation. Perceptual quality is assessed using no-reference metrics including NIQE (Natural Image Quality Evaluator), BRISQUE (Blind/Referenceless Image Spatial Quality Evaluator), and HyperIQA su2020blindly . For audio-visual synchronization, we measure LMD (Landmark Distance) between predicted and ground truth facial landmarks in the mouth region, and LSE-C (Lip Sync Error - Confidence) to evaluate lip synchronization quality.
For the AIGC-LipSync benchmark, we report the Generation Success Rate across all 615 videos and specifically for stylized characters. This metric shows the percentage of videos that are successfully synchronized and pass human verification. This evaluation is essential for universal lip synchronization in AI-generated content, where traditional metrics may not fully capture the challenges of stylized characters, extreme poses, and other atypical visual conditions.
[…]
Figure 3: Qualitative comparison with previous methods across diverse subjects and phonemes. Our approach produces more accurate lip synchronization and better identity preservation.
Table 3: User study results comparing various audio-driven lip-sync methods.
Metric Wav2Lip
Video
ReTalking
TalkLip IP-LAP Diff2Lip MuseTalk
Latent
Sync
Ours
Lip Sync Accuracy 2.684 2.769 1.940 2.359 3.000 2.632 3.812 3.923
Identity Preservation 2.410 2.786 2.222 2.991 2.889 3.197 3.658 4.128
Timing Stability 2.556 2.376 2.162 3.034 2.812 3.145 3.581 4.043
Image Quality 2.017 2.607 1.889 3.171 2.402 3.094 3.632 4.051
Video Realism 2.120 2.419 1.838 2.316 2.479 2.761 3.453 3.872
4.3 Qualitative Evaluation
We present qualitative comparisons between OmniSync and existing methods in Fig. 3. Our approach produces more natural facial expressions and superior lip synchronization. Due to lip shape leakage, IP-LAP zhong2023identity and LatentSync li2024latentsync frequently fail at mouth shape modification, resulting in poor lip synchronization effects. MuseTalk zhang2024musetalk and VideoReTalking cheng2022videoretalking modify lip movements but frequently lose identity and visual quality. Diff2Lip mukhopadhyay2024diff2lip and Wav2Lip prajwal2020lip commonly exhibit lip sync errors, mouth artifacts, and identity drift, particularly in challenging or stylized cases. In contrast, OmniSync consistently maintains identity details and generates realistic, expressive lip movements, demonstrating robust performance. Our approach effectively balances audio and visual cues, addressing the challenge of weak audio conditioning.
User Study. To assess perceptual quality, we conducted a user study with 39 participants evaluating 32 video sets generated by OmniSync and seven competing methods, with a standardized Cronbach’s $\alpha$ coefficient of 0.98. Participants rated each video on a 5-point Likert scale across five criteria: Lip Sync Accuracy, Character Identity preservation, Timing Stability, Image Quality, and Video Realism. As shown in Tab. 3, OmniSync outperformed all competitors across all metrics, achieving superior scores in Lip Sync Accuracy (3.923 vs. 3.812 for LatentSync), Character Identity (4.128 vs. 3.658), Timing Stability (4.043 vs. 3.581), Image Quality (4.051 vs. 3.632), and Video Realism (3.872 vs. 3.453). These results confirm OmniSync’s superior ability to generate high-quality talking videos.
4.4 Ablation Study
To clarify the contributions of each core component in our framework, we conduct an ablation study targeting three key modules: the timestep-dependent data sampling strategy, progressive noise initialization, and the Dynamic Spatiotemporal Classifier-Free Guidance (DS-CFG) mechanism. Quantitative results are presented in Tab. 4, and corresponding visual examples are shown in Fig. 4.
[…]
We also compare our proposed DS-CFG with both low and high static CFG settings. As illustrated in Fig. 4, low CFG provides insufficient audio conditioning, resulting in under-articulated lip movements (LSE-C: 4.16), whereas high CFG improves synchronization (LSE-C: 7.10) but introduces noticeable artifacts and distortions in facial details. In contrast, DS-CFG achieves an optimal balance by applying strong localized guidance in early diffusion stages and gradually reducing it in later steps. These results confirm that dynamic control across temporal and spatial dimensions is essential for producing expressive and visually coherent lip synchronization in generative video content.
Table 4: Ablation study for our method.
Methods FID ↓ FVD ↓ CSIM ↑ NIQE ↓ BRISQUE ↓ HyperIQA ↑ LSE-C ↑
Ours 15.710 287.168 0.814 5.321 29.588 57.288 7.06
w/o Timestep-Dependent
Sampling Strategy
21.552 549.768 0.727 5.462 30.346 56.204 7.00
w/o Progressive Noise
Initialization
16.731 361.282 0.805 5.349 29.789 56.511 7.03
w/ Low Static CFG - - - 5.359 29.724 56.568 4.16
w/ High Static CFG 22.725 348.335 0.782 5.473 29.678 56.289 7.10
Figure 4: Ablation study for timestep-dependent sampling strategy and different CFG settings.
5 Conclusion
In this paper, we introduce OmniSync, a universal lip synchronization framework for diverse content that addresses critical limitations of traditional approaches. Our three key innovations—a mask-free training paradigm eliminating mask dependencies, a flow-matching-based progressive noise initialization strategy ensuring identity preservation, and dynamic spatiotemporal Classifier-Free Guidance balancing synchronization with visual quality—collectively enable precise lip movements across diverse visual representations. To support systematic evaluation in this field, we establish the AIGC-LipSync Benchmark, the first comprehensive framework for assessing lip synchronization in varied AIGC contexts. Extensive experiments demonstrate OmniSync’s superior performance across challenging scenarios, establishing a robust foundation for integrating precise lip synchronization into the broader AI video generation ecosystem.
Appendix A Training Procedure Details
A critical aspect of our training methodology is the timestep-dependent sampling strategy. The core idea behind this strategy is to provide the model with different types of training data according to the different denoising stages of the diffusion model. This optimizes learning efficiency and final generation quality, especially for the specific task of lip synchronization.
Considerations for Early Denoising Stages and Link to Progressive Noise Initialization
The primary requirement of the lip synchronization task is to modify only the lip region to match new audio while preserving the head pose, identity features, and background environment of the input video. The early stages of a traditional diffusion process starting from pure random noise (i.e., when the noise level is very high, and $t$ approaches the total number of steps $T$) are mainly responsible for generating the macroscopic structure of the image, including pose and general identity outlines. However, for lip synchronization, the input video itself already provides this macroscopic structural information. Since our goal is to not change the pose and background, learning this already-given information from scratch (starting from random noise) is not only inefficient but can also introduce unnecessary variability, potentially leading to discrepancies in pose or identity between the final generated video and the original.
Therefore, during inference, we employ the Progressive Noise Initialization strategy. This involves directly adding a controlled level of noise (corresponding to a high starting timestep $t_{start}$, e.g., $\tau=0.92$) to the original video frames (which we refer to as "base frames") and then proceeding with the subsequent denoising steps from this $t_{start}$. This is equivalent to "skipping" the early denoising process from pure noise to $t_{start}$. The rationale is that the base video already contains all the macroscopic structural information (pose, identity, background) that we wish to preserve. By directly noising the base video, we provide the model with an initial state that already possesses the correct pose and identity, allowing it to focus the subsequent denoising process more on the precise editing of the lip region.
Timestep-Dependent Sampling Strategy in Training
To enable the model during the training phase to adapt to this inference mode of "skipping early structure generation" and to effectively learn the specific knowledge required for lip synchronization, we designed the following timestep-dependent sampling strategy:
  • For higher noise levels (early diffusion stages, e.g., $t\textgreater 850$): At this stage, even when starting from random noise, the model is primarily learning to construct basic facial structures and poses. To provide the model with the most stable and relevant learning signals, we sample from pseudo-paired data. We specifically selected the MEAD dataset as our source of pseudo-paired data. The MEAD dataset, recorded under controlled laboratory conditions, offers several crucial advantages:
    • Fixed Recording Conditions: Multiple emotional expressions from the same subjects were captured with fixed camera positions and consistent lighting. This results in video sequences where the facial structure and identity remain constant, with variations only in lip shapes and expressions.
    • Natural Pseudo-Pairs: Within the same identity, frames from different utterances form natural pseudo-pairs—they maintain nearly identical head poses and environmental conditions but differ in lip configurations.
    • Multi-View Capture: The multi-view setup in MEAD further enriches our training data by providing consistent identity representations across different angles. This enables the model to learn pose-invariant facial structures more robustly during the early diffusion timesteps.
    By using this carefully curated pseudo-paired data, the model, in the early (high-noise) stages, learns how to begin constructing facial features under given (or nearly given) pose and identity conditions, laying a solid foundation for subsequent lip generation. At this point, the model needs to learn how to recover a structure from the noise that is highly consistent in pose and identity with the input condition ($V_{cd}$, the noised version of the base video), while preparing for the generation of the target lip shape (the noised version of $V_{ab}$). Therefore, a strong correspondence in pose between the input ($x_{t}$ from noised $V_{cd}$) and the target (noised $V_{ab}$) is critical.
  • For lower noise levels (middle and late diffusion stages, e.g., $t\leq 850$): As the diffusion process enters the middle and late stages, the noise level gradually decreases. At this point, by visualizing $x_{t}$ at different stages (similar to the analysis in works like VideoJAM chefer2025videojam ), we can observe that the main contour information, facial structure, and identity features have become relatively clear. In other words, the model has already "seen" the character’s pose and general identity.
    • Shift in Model’s Learning Focus: The primary task of the model at this stage is no longer to construct macroscopic structures but to finely sculpt the lip shape and texture details according to the audio signal, ensuring a natural blend with the rest of the face.
    • Rationale for Using Unpaired Data: Since macroscopic information such as pose and identity is largely established (guaranteed by early-stage learning and/or progressive initialization at inference), the model no longer heavily relies on strict pose pairing between input and target. Therefore, we can transition to sampling from our broader and more diverse collection of arbitrary/unpaired videos (from the YouTube dataset). While this data may not match any specific input video $V_{cd}$ in terms of pose, identity, or scene, it provides a vast number of lip movement samples under different speaking contents and styles. The model at this stage learns a more generalized "audio-to-lip-shape" mapping and how to apply this mapping to an $x_{t}$ that already possesses basic contours, to generate lip shapes synchronized with the target audio $A_{ab}$, and to refine local details and realism. This strategy significantly expands the diversity of training data, enhancing the model’s generalization capabilities to handle various real-world lip synchronization scenarios.
In summary, our timestep-dependent sampling strategy, combined with progressive noise initialization at inference, allows OmniSync to efficiently focus on the core challenges of lip synchronization. By using pseudo-paired data in the early stages to stabilize structural learning and leveraging large-scale unpaired data in the middle and late stages to learn diverse lip expressions, the model ensures both identity and pose consistency in the generated videos and possesses strong lip generation and generalization capabilities.
Appendix B Comparison with Portrait Animation Methods
Our OmniSync framework presents distinct advantages when compared to state-of-the-art Portrait Animation methodologies, such as EMO tian2024emo , EchoMimic chen2025echomimic , Hallo xu2024hallo , and Sonic ji2024sonic . While these models can achieve good lip synchronization and generalize across diverse visual styles, they often struggle to fully preserve the unique identity, texture, and dynamic characteristics of source video material. This can lead to animations that, despite accurate lip sync, may appear somewhat generic or lose fine-grained subject resemblance.
In contrast, OmniSync’s training paradigm, formulated as $(V_{cd},A_{ab})\mapsto V_{ab}$, utilizes multi-frame video input ($V_{cd}$) instead of a single static image. This video input inherently encodes richer information, including subtle facial dynamics and individual speaking styles over a temporal window, beyond static texture details. Conceptually, while the input $V_{cd}$ in our framework serves a conditioning role analogous to the single image in portrait animation approaches, the crucial difference lies in the temporal dimension of $V_{cd}$, which allows for the preservation of more stylistic and identity-specific information.
This architectural choice enables OmniSync to maintain strong generalization capabilities while achieving high fidelity to the source. As shown in Fig. 5, our qualitative comparison indicates that methods such as EchoMimic chen2025echomimic , Hallo3 cui2024hallo3 , and Sonic ji2024sonic face challenges in maintaining identity, which may result in less realistic outcomes, while OmniSync continuously generates synchronized speech, better preserving the distinctive texture quality and inherent speaking style of the input video. This holds true even when applied to out-of-distribution subjects, such as non-human or highly stylized characters. Our model effectively learns to transfer primarily the lip movements dictated by the target audio, while retaining other visual characteristics of the source video segment. Consequently, OmniSync’s outputs appear as more natural integrations with the original content.
Figure 5: Comparison with portrait animation methods. Visual comparison between our OmniSync framework and other approaches (EchoMimic, Hallo3, and Sonic), demonstrating our method’s superior ability to preserve identity and natural speaking style while maintaining accurate lip synchronization.
Appendix C Lip Shape Leakage in Lip Synchronization
Lip shape leakage occurs when original lip movements from source videos persist in the synchronized output despite attempts to replace them with new audio-driven movements. This phenomenon commonly affects traditional lip synchronization methods that rely on masked-frame inpainting. When using explicit masks to isolate mouth regions, the boundaries between what should be modified and preserved often become ambiguous, allowing original lip shapes to "leak" into the generated content. Additionally, the relatively weak conditioning signal provided by audio compared to visual information exacerbates this problem, as models may default to preserving visual aspects of the original video rather than fully replacing them.
The impact of lip shape leakage is significant for synchronization quality. Viewers perceive temporal inconsistencies where mouth movements partially match both the original video and target audio, creating unnatural motion patterns. Critical phonemes requiring distinct mouth shapes (such as bilabial or labiodental sounds) may not form correctly, reducing visual intelligibility and breaking the illusion of natural speech. As demonstrated in Fig. 6, methods like IP-LAP and LatentSync frequently show this limitation, particularly for challenging phonemes that require distinct mouth formations.
Figure 6: Comparison of lip synchronization methods showing how our approach effectively prevents lip shape leakage from the original video, while competing methods struggle with this issue (highlighted in red for IP-LAP and LatentSync).
OmniSync addresses lip shape leakage through its mask-free training paradigm and dynamic spatiotemporal guidance mechanism. Rather than using explicit masks with hard boundaries, our diffusion-based approach directly modifies frames based on audio conditioning, allowing the model to determine appropriate modification regions dynamically. Our DS-CFG further mitigates leakage by applying stronger guidance around mouth regions, ensuring audio conditioning overcomes any tendency for original lip shapes to persist in the output. These innovations effectively eliminate lip shape leakage, enabling precise and consistent lip synchronization across diverse visual representations, including stylized characters where traditional face alignment techniques often fail.
Figure 7: Comparison of lip synchronization without (left) and with (right) Video Description conditioning. Text guidance produces more pronounced lip movements with clearer dental visibility across diverse subjects, as shown in the magnified mouth regions.
Appendix D Impact of Video Description on Lip Clarity and Movement Amplitude
We conducted an ablation study to evaluate the effect of video description conditioning on lip synchronization quality. As shown in Fig. 7, we compare results generated without video description (left column) to those with video description conditioning (right column). The qualitative comparison clearly demonstrates the impact of textual guidance on lip movement characteristics.
During training, we labeled videos with descriptive prompts such as "A person speaking loudly with clear facial and tooth movements" to establish associations between textual descriptions and specific lip characteristics. This approach allows the model to learn correlations between descriptive language and visual speech attributes. At inference time, these text descriptions serve as an interpretable control mechanism, enabling adjustment of lip clarity and movement amplitude through prompt engineering without requiring model retraining. The examples demonstrate how this text-guided approach results in more expressive and visually distinct lip synchronization, enhancing overall perceptual quality for viewers.
From the visual results, two key improvements are immediately apparent with video description conditioning. First, the lip movements exhibit greater amplitude and expressiveness, particularly evident in the middle and bottom rows where the subjects display more pronounced mouth openings. Second, there is notably improved clarity of dental structures, with teeth being more visible and defined in the right column examples. This enhancement in visual detail contributes significantly to the realism and comprehensibility of the generated speech.
Appendix E Benchmark Construction Details
To comprehensively evaluate the performance of lip synchronization methods within the current AI-Generated Content (AIGC) environment, we have meticulously constructed the AIGC-LipSync Benchmark. This benchmark comprises a total of 615 video clips, all generated by leading text-to-video (T2V) or image-to-video (I2V) models. These videos primarily originate from advanced models including Kling, Dreamina, Wan wang2025wan , and Hunyuan kong2024hunyuanvideo , with all raw video materials downloaded from publicly accessible communities such as Civitai, ensuring a diverse and representative collection.
Figure 8: AIGC-LipSync Benchmark Examples (I): Showcasing non-human characters and landscape-oriented video materials included in the benchmark. This highlights its coverage of diverse subject types and video formats, designed to test the model’s lip synchronization capabilities on non-traditional visual inputs.
Figure 9: AIGC-LipSync Benchmark Examples (II): Presenting diverse male and female character materials from the benchmark. These are used to evaluate the model’s lip synchronization effectiveness and identity preservation across various human figures, particularly in the rendition of facial features and nuanced expressions.
This benchmark has been specifically curated to include a variety of AI-generated content that poses significant challenges for lip synchronization. Among all data, there are approximately 450 videos featuring realistic human subjects, around 125 videos of stylized characters with distinct artistic styles, and a smaller set of approximately 40 videos depicting more challenging, atypical humanoid characters. This composition is designed to span a wide spectrum of visual representations, from photorealistic to highly abstract, thereby enabling a more rigorous test of model generalization capabilities and robustness. These video clips have an average duration of approximately 6 seconds, an average resolution of 976x1409 pixels, and an average frame rate maintained at 30.00 FPS, providing sufficient data for detailed synchronization analysis. As illustrated in Fig. 8 and 9, representative examples from our benchmark showcase its content diversity and the inherent challenges it presents.
Appendix F Ethical Considerations
Our OmniSync framework, while advancing the field of lip synchronization, raises important ethical considerations regarding potential misuse. As with any technology capable of manipulating facial content, there exists risk for creating misleading or deceptive media that could contribute to misinformation or deepfakes. We acknowledge this responsibility and have intentionally focused our research on improving existing video content rather than enabling impersonation or fabrication of speech.
To mitigate potential harms, we recommend implementing watermarking or content provenance solutions when deploying this technology commercially. Additionally, the creation of the AIGC-LipSync Benchmark emphasizes evaluating performance on stylized characters and AI-generated content, steering applications toward creative and entertainment purposes rather than realistic human impersonation.
We are committed to transparency regarding the capabilities and limitations of our approach. The technical advancements presented in OmniSync are published to advance scientific understanding and enable beneficial applications in areas such as film production, accessibility services, and educational content. We encourage the research community to continue developing detection methods for synthetically modified content alongside improvements in generation quality.
Appendix G Acknowledgments
We would like to express our sincere gratitude to the vibrant community at Civitai for their generous sharing of creative content. The majority of videos in our AIGC-LipSync Benchmark were collected from this platform, showcasing the diverse and artistic AI-generated content that creators within our community continuously produce and share. Their contributions have been instrumental in enabling comprehensive evaluation of lip synchronization methods across the full spectrum of modern AI-generated visual content.